OpenStudy (anonymous):

Pat Maninan, a customer of Brown Co., will pay $400 for a new kitchen table. Brown has a 55% markup on selling price. What is the most Brown can pay for this kitchen table?

OpenStudy (anonymous):

this is my last question!

OpenStudy (ash2326):

We all are here to help. Ask as many as you want:) Welcome to Open Study:D

OpenStudy (dumbcow):

4000 is selling price Let p be cost. because of 55% markup markup/4000 = 0.55 markup = .55 * 4000 = 2200 p = 4000 - 2200

OpenStudy (anonymous):

so it wud be 1800

OpenStudy (ash2326):

@dumbcow it's asking for what max he could pay. I think $400 is not selling price

OpenStudy (dumbcow):

oh oops i put 4,000 and how else would you determine selling price though?

OpenStudy (ash2326):

Markup= 55% of 400=220 Selling Price= $400+$220=$620

OpenStudy (ash2326):

The most Brown can pay for this kitchen table is $620

OpenStudy (anonymous):

so then it wud be 620?

OpenStudy (ash2326):

Yeah I think so

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

Brown is the seller, not the buyer

OpenStudy (anonymous):

i have 2 more questions to do!

OpenStudy (anonymous):

with only 35 minutes to do them

OpenStudy (ash2326):

Oh you are right:) Let the cost be x 55% markup Markup= 0.55x so 1.55x=400 so x=258.06 so the most Brown can pay is $258.06

OpenStudy (anonymous):

thnk u(:

OpenStudy (kropot72):

The most Brown can pay is $180 in order to make his required mark up on the selling price of $400.

OpenStudy (anonymous):

ok now im confused >.<

OpenStudy (dumbcow):

the markup is 55% of the selling price (400) --> markup = 220 cost = 400 - 220 = 180

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

ash, you were setting markup as 55% of cost price